### v3

@@ 1 Introduction
In contrast, pre-trained MLFFs, that make use of extensive DFT datasets, offer explicit relaxation of adsorbates on catalyst surfaces and a significant speed-up (a factor of 104 or more) compared to DFT calculations while maintaining quantum mechanical accuracy [28, 29] making them suitable for high-throughput catalyst discovery workflows. Modern Sabatier principle-based approaches also utilize MLFFs or specialized models to find (global) minimum adsorption energies across multiple material facets [21, 13] providing a scalable and transferable framework for screening a broad range of catalysts. Although several studies emphasize the importance of various catalyst facets [9, 30], these approaches predominantly utilize data from individual facets for material characterization. Consequently, the challenge persists: how can we effectively predict catalytic performance without limiting our scope to specific material families or facet orientations?
@@
Secondly, we aim to establish a high-throughput and ML-enhanced workflow to accelerate the screening of catalytic materials using our newly formulated descriptor. Traditional density functional theory (DFT) approaches are computationally prohibitive for large-scale studies. To overcome this limitation, we leverage MLFFs from the Open Catalyst Project (OCP) [33, 34], enabling the rapid and accurate computation of adsorption energies. Our workflow generates an extensive dataset of AEDs, capturing over 877,000 adsorption energies for nearly 160 materials relevant to the CO2 to methanol conversion reaction. To target the enhanced reliability of our prediction and the effective use of ML models in catalyst discovery, we design a robust validation protocol.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [30] highlighted the presence of surface-bound radicals such as *H (hydrogen atom), *OH (hydroxy group), *OCHO (formate), and *OCH3 (methoxy) as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected these adsorbates for our AEDs calculations. Please note that the notation for formate can vary across the literature, e.g., *HCOO [30, 39] and HCOO* [40]. With the help of fairchem repository tools by OCP [41], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that six materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
The validation step is connected with the final data-cleaning when we exclude any material with an EMAE surpassing the threshold of 0.25 eV. Consequently, 29 materials were expunged from our dataset, retaining 159 materials. Most materials flagged for significant EMAEs exhibited magnetic properties, exemplified by materials like MnCo, MnGa, or FeCo. Magnetism presents significant challenges for the non-spin-polarized DFT calculations used in OC20 and in this work. A complete list of estimated MAEs for the remaining 159 materials is accessible in [43].
@@
Inspection of Fig. 3 and Fig S1 reveals that adsorption energies span a wide interval from -7.42 eV to 2.40 eV. We included energies above zero, although positive adsorption energies are typically indicative of molecular desorption. However, the adsorption energies reported in this work do not include entropy and pressure terms, which could shift the energies to more negative values. Secondly, the adsorption energy of radicals is somewhat ill-defined, if different desorption channels are conceivable. Since our objective is to achieve a qualitative comparison across materials, the price energy zero is of no relevance, as long as it is chosen consistently.
@@
Figure 3: AEDs for selected subset of materials: (a) AEDs for 26 material with the amplitude of the distribution encoded in the intensity of the corresponding color; detailed AEDs for (b) K, (c) Y3In5, (d) Ni, (e) CuZn, (f) NiZn, (g) ZnRh. The numbers on the y-axis of panel (a) indicate the cluster numbers defined in Fig. 4.
@@ 3.5 Unsupervised Learning
Table S1 – All 159 materials used in final evaluation; Table S2 – All materials removed from the final evaluation; Figure S1 – Adsorption energy distributions and hierarchical clustering of all 159 materials; Table S3 – Statistical analysis on the AEDs for a subset of materials shown in Fig. 3.

### v5

@@ 1 Introduction
In contrast, pre-trained MLFFs, that make use of extensive DFT datasets, offer explicit relaxation of adsorbates on catalyst surfaces and a significant speed-up (a factor of 104 or more) compared to DFT calculations while maintaining quantum mechanical accuracy [28, 29], making them suitable for high-throughput catalyst discovery workflows. Modern Sabatier principle-based approaches also utilize MLFFs or specialized models to find (global) minimum adsorption energies across multiple material facets [21, 13] providing a scalable and transferable framework for screening a broad range of catalysts. Although several studies emphasize the importance of various catalyst facets [9, 30], these approaches predominantly utilize data from individual facets for material characterization. Consequently, the challenge persists: how can we effectively predict catalytic performance without limiting our scope to specific material families or facet orientations?
@@
Secondly, we aim to establish a high-throughput and ML-enhanced workflow to accelerate the screening of catalytic materials using our newly formulated descriptor. Traditional density functional theory (DFT) approaches are computationally prohibitive for large-scale studies. To overcome this limitation, we leverage MLFFs from the Open Catalyst Project (OCP) [33, 34], enabling the rapid and accurate computation of adsorption energies. Our workflow generates an extensive dataset of AEDs, capturing over 877,000 adsorption energies for nearly 160 materials relevant to the CO2 to methanol conversion reaction. To target the enhanced reliability of our predictions and the effective use of ML models in catalyst discovery, we design a robust validation protocol.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [30] highlighted the presence of surface-bound radicals such as *H (hydrogen atom), *OH (hydroxy group), *OCHO (formate), and *OCH3 (methoxy) as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected these adsorbates for our AEDs calculations. Please note that the notation for formate can vary across the literature, e.g., *HCOO [30, 39] and HCOO* [40]. With the help of fairchem repository tools by OCP [41], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that seven materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
The validation step is connected with the final data-cleaning when we exclude any material with an EMAE surpassing the threshold of 0.25 eV. Consequently, 29 materials were expunged from our dataset, retaining 158 materials. Most materials flagged for significant EMAEs exhibited magnetic properties, exemplified by materials like MnCo, MnGa, or FeCo. Magnetism presents significant challenges for the non-spin-polarized DFT calculations used in OC20 and in this work. A complete list of estimated MAEs for the remaining 158 materials is accessible in [43].
@@
Inspection of Fig. 3 and Fig S1 reveals that adsorption energies span a wide interval from -7.42 eV to 2.40 eV. We included energies above zero, although positive adsorption energies are typically indicative of molecular desorption. However, the adsorption energies reported in this work do not include entropy and pressure terms, which could shift the energies to more negative values. Secondly, the adsorption energy of radicals is somewhat ill-defined if different desorption channels are conceivable. Since our objective is to achieve a qualitative comparison across materials, the price energy zero is of no relevance, as long as it is chosen consistently.
@@
Figure 3: AEDs for selected subset of materials: (a) AEDs for 25 materials with the amplitude of the distribution encoded in the intensity of the corresponding color; detailed AEDs for (b) K, (c) Y3In5, (d) Ni, (e) CuZn, (f) NiZn, (g) ZnRh. The numbers on the y-axis of panel (a) indicate the cluster numbers defined in Fig. 4.
@@ 3.5 Unsupervised Learning
Table S1 – All 158 materials used in final evaluation; Table S2 – All materials removed from the final evaluation; Figure S1 – Adsorption energy distributions and hierarchical clustering of all 158 materials; Table S3 – Statistical analysis on the AEDs for a subset of materials shown in Fig. 3.